Determine the p-value given the stated hypothesis and test statistic value (Z) Ho: μ = 120 H1: μ ≠ 120; z = 1.92
Answer #1

Given that, the null and alternative hypotheses are,

Ho: μ = 120

H1: μ ≠ 120

Test statistic is, z = 1.92

since, it is two-tailed test,

p-value = 2 * P(Z > 1.92) = 2 * 0.0274 = 0.0548

p-value = 0.0548

Note: P(z > 1.92) = 1 - P(Z < 1.92) = 1 - 0.9726 = 0.0274